OpenStudy (anonymous):

What is the solution to the equation? c + (–2) = –0.7

OpenStudy (briannaontimeforreal):

ok i don't get it tell me where you are stuck

OpenStudy (anonymous):

i just need the answer for c

OpenStudy (anonymous):

to get c + (–2) = –0.7 -7=7/10 c-2==-7/10 I think that would be c= 1.3 0r c-1.3=0

OpenStudy (anonymous):

ok thanks i guess lol

OpenStudy (anonymous):

hope you understood :)

OpenStudy (anonymous):

yeah

OpenStudy (briannaontimeforreal):

1.3